A blue jay has a nest with 7 eggs. There is a 50% chance that they will hatch to be male.


What is the probability that the nest will hatch 3 males?


What is the probability that the nest will hatch more than 3 males?


What is the probability that the nest will hatch at most 3 males?
Answer #1

Since Blue jay has a nest with 7 eggs and there is a 50% chance that they will hatch to be male.

Number of trials = n = 7

And probability of egg will hatch male = p = 0.5

(a) the probability that the nest will hatch 3 males

Number of success = X = 3

Applying the binomial probability formula

p(X) = \binom{n}{X}*p^{X}*(1-p)^{n-X}

Where

\binom{n}{X} = \frac{n!}{X!(n-X)!}

Therefore

p(X) =\frac{n!}{X!(n-X)!}*p^{X}*(1-p)^{n-X}

p(3) =\frac{7!}{3!(7-3)!}*0.5^{3}*(1-0.5)^{7-3}​​​​​​

p(3) =0.2734

So the probability that the nest will hatch 3 males is 0.2734

(b)  the probability that the nest will hatch more than 3 males

p(X>3) = p(4) + p(5) +p(6) +p(7)

So

p(4) =\frac{7!}{4!(7-4)!}*0.5^{4}*(1-0.5)^{7-4} = 0.2734

p(5) =\frac{7!}{5!(7-5)!}*0.5^{5}*(1-0.5)^{7-5} = 0.1641

p(6) =\frac{7!}{6!(7-6)!}*0.5^{6}*(1-0.5)^{7-6} = 0.0547

p(7) =\frac{7!}{7!(7-7)!}*0.5^{7}*(1-0.5)^{7-7} = 0.0078

p(X > 3) = 0.2734+0.1641+0.0547+0.0078 = 0.5​​​​​​

So the probability that the nest will hatch more than 3 males is 0.5

(c)the probability that the nest will hatch at most 3 males

p(X\leq 3) = p(3)+p(2)+p(1)+p(0)

0r,

p(X\leq 3) = 1-p(X> 3)

Since we have calculated p(X>3) in last part

p(X\leq 3) = 1-0.5 = 0.5

So the probability that the nest will hatch at most 3 males is 0.5.